OpenStudy (anonymous):

5x+9y+4=0 has the x intercept -4/5 and the y intercept -4/9 It defines a straight line of slope ???????

OpenStudy (anonymous):

Put the equation in slope-intercept from (meaning solve it for y). The slope is then the coefficient of the x term.

OpenStudy (anonymous):

Sorry, previous post should have said slope-intercept "form", not "from"

OpenStudy (anonymous):

is slope mx+b

OpenStudy (anonymous):

slope is m. b is the y-intercept

OpenStudy (anonymous):

ok so how would i set this up as a formula

OpenStudy (anonymous):

If I understand your question correctly, you just need to re-arrange the equation you already have to put it in slope-intercept form. First, you need to eliminate any terms that don't include "y" from the left side of the equation, so you would subtract "5x" and "4" from both sides. Second, You would divide both sides by 9 to eliminate the 9 on the left. That would put your equation in slope-intercept form.

OpenStudy (anonymous):

At that point, you can identify "m", which is the slope of your line. "m" is the number or fraction in front of the x term when the equation is in slope-intercept form.

OpenStudy (anonymous):

ok so the equation right now should be y= -5x-4

OpenStudy (anonymous):

After subtraction, you would have: 9y=-5x-4 Then, after division, you would have: y=-5/9 x - 4/9

OpenStudy (anonymous):

ok so the slope would be -5/9?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok thanks that makes sense

OpenStudy (anonymous):

5x+9y+4=0 \[y=(-5x-4)/9\] \[y= (-5/9)x - 4/9\] m= -5/9 b = -4/9